# A New Approach to Solving SMAC Task: Generating Decision Tree Code from Large Language Models

### v1

@@ 4 Experiment
In this section, we evaluate our proposed framework on a variety of scenarios from the SMAC micro-management tasks, which require the use of various individual and combined skills. The results demonstrate that through the planning, coding, and critic phases, the LLMs are able to select appropriate skills, generate Python code that complies with python-sc2 requirements, and successfully win the combat scenarios. The code is available at https://github.com/devindeng94/LLM-SMAC.
4.1 Environmental Settings
From the 23 SMAC task scenarios, we selected 14 maps that represent a range of basic skill applications and complex skill combinations. These maps are:
  1. 1.
    3m, 2s3z, and 8m, which emphasize ’focus firing’.
  2. 2.
    2m_vs_1z, 3s_vs_3z, 3s_vs_4z, and 3s_vs_5z, which require ’hit-and-run’ tactics.
  3. 3.
    2c_vs_64zg and corridor, where leveraging ’terrain’ advantages is key.
  4. 4.
    MMM, 2s_vs_1sc, and 3s5z, where ’health and shield management’ is the primary focus.
  5. 5.
    1c3s5z and bane_vs_bane, which emphasize ’kiting’ and maximizing damage output.
While we describe the primary skills required for each map, auxiliary skills are also necessary to execute successful strategies. Additionally, we set the adversary difficulty to ’Very Hard’, which corresponds to level 7 in the original SMAC MARL tasks. We validate the generated code within the StarCraft II client using 10 different random seeds. Both the time taken for planning and coding are recorded as part of the results.
For the sake of convenience, we use the open-source DeepSeek-V2.5 as the base model for the planner, coder, and critic modules. However, these models can be replaced with more advanced alternatives, such as GPT-4.
@@ 4.2 Main Results
Map Name WR % Plan Code Planned Strategy
3m 100% 1 2 Focus Fire
8m 100% 1 5 Focus Fire
10m_vs_11m 60% 3 19 Focus Fire, Grouping
2s3z 100% 7 31 Focus Fire, Grouping
3s5z 100% 4 22 Focus Fire, Grouping
2m_vs_1z 100% 1 4 Hit and Run
3s_vs_3z 100% 1 3 Focus Fire, Hit and Run
3s_vs_4z 100% 2 14 Focus Fire, Hit and Run
3s_vs_5z 100% 2 14 Focus Fire, Hit and Run
2c_vs_64zg 100% 4 18 Health and Shield Management, Terrain Advantage
corridor 100% 3 17 Health and Shield Management, Terrain Advantage
MMM 100% 2 15 Focus Fire, Hit and Run, Spread out and Heal
2s_vs_1sc 100% 1 23 Health and Shield Management, Terrain Advantage
bane_vs_bane 100% 3 14 Spread
1c3s5z 100% 2 11 Engagement, Hit and Run, Flanking
Table 1: The winning rate over 10 times validation, the strategy planning rounds, the code generation rounds, and the planning results of each map in SMAC.
We present the results of win rates, planning rounds, coding rounds, and the final strategies planned by the LLM for each map in Table 1. As shown in the table, the LLM-generated code achieves near-perfect win rates on easier tasks and competitive results on more difficult tasks. Leveraging the LLM’s prior knowledge, the planner is able to select optimal strategies with relative ease, and thanks to the high coding capabilities of the DeepSeek model, most correct Python scripts are generated within 30 rounds of iteration.
@@
The planner LLM outlines the strategy framework by combining skills and defining the conditions under which they should be used. According to the results, ’Focus Fire’ emerges as one of the most common and effective skills across SMAC maps, as it quickly reduces the number of enemy units. For many of the easier tasks, once the planner identifies and selects the ’Focus Fire’ skill, the likelihood of generating a winning code script significantly increases. As a result, planning rounds for maps requiring focus fire tend to be much shorter. Conversely, on maps where focus fire is less effective, such as the corridor map, the LLM requires more rounds to produce alternative skills after receiving feedback from the rollout results.
@@
Figure 3: Left An example planner output that provides three potential skills: focus fire, hit and run, and spread out and heal for the medivac unit. The planner also describes how to use each skill and when to use the skill. Right A snapshot of the MMM rollout process. According to the planner strategy, the Marines and Marauders are attacking, two Marauders are spreading out to take damage, two Marines are Running back for healing, and the Medivac is healing a Marine unit.
@@
Figure 4: An example coder output from the first round and the code after the critic promotion round. Both of the two example codes follow the strategy but the latter code is more precise.
As shown in Figure 4, starting from a strategy skeleton, we present the coding result from the first round and the refined code following the promotion prompt from the critic module. During the rollout process, the initial code (on the left) achieves 4 wins out of 10 simulations, while the refined code (on the right) achieves a perfect 10 out of 10 wins. Compared to the original code, the improved version refines retreat distances and adjusts the health and shield threshold values. Additionally, the refined code includes group coordination logic, further boosting the win rate.
@@
Beyond value adjustments and skill additions, the coder can also enhance the implementation by accounting for more situational factors. As illustrated in Figure 5, in the kiting strategy, agents should retreat in the opposite direction to maintain a minimum distance from the closest enemy. However, on smaller maps, units may get stuck at the map’s edge, resulting in the retreating behavior being taken but failing. After several rounds of in-context learning improvements, the critic identifies that the planned strategy is sound but the implementation lacks detail. Consequently, the coder begins dynamically calculating the retreat direction in the kiting method, improving the strategy’s overall effectiveness.
Figure 5: An example coder output from the first round and after the critic promotion round given a main strategy skeleton. Both of the two example codes follow the strategy but the latter code is more precise.
4.3 Discussion
@@
Figure 7: An example indicates that the strategy learned from the 3s_vs_4z scenario is that one of the units is responsible for kiting and another two units are responsible for attacking. Because the universal strategy is learned on the homogeneous scenario, the strategy is also suitable for 3, 4, and 5 enemy Zealots.
@@
Vanilla reinforcement learning models often suffer from low exploration efficiency, as they must train from scratch. Although some approaches in multi-agent reinforcement learning emphasize coordinated exploration techniques, these methods still explore at the action level, which limits their efficiency. In contrast, our LLM-SMAC framework leverages the prior knowledge of LLMs to explore at the skill level in the planner phase and at the decision-tree implementation level in the coder phase. This significantly enhances efficiency, allowing us to solve hard SMAC tasks in just tens of rounds of planning and code generation, whereas MARL algorithms typically require millions of timesteps over tens of thousands of iterations.
@@
Map descriptions provide crucial terrain advantages that enrich strategic planning. Some units, like the Colossus, can use the terrain to their advantage, such as climbing cliffs to kite enemies. Other units may use map edges to block choke points, reducing the risk of being surrounded by Zerglings. As illustrated in Figure 8, two Colossi can exploit the terrain by climbing up and down a cliff, while most enemy Zerglings remain on the other side. Without this map information, strategies might unintentionally rely on map edge advantages. Another example is the corridor scenario, where units should hold choke points to prevent being surrounded by Zerglings. Without the choke point, the generated code is hard to win the corridor combat.

### v2

@@ 4 Experiment
In this section, we evaluate our proposed framework on 23 scenarios from the SMAC micro-management tasks, which require the use of various individual and combined skills. We also test our SMAC-R1 on 10 newly-designed similar scenarios. The results demonstrate that through the planning, coding, and critic phases, the LLMs are able to select appropriate skills, generate Python code that complies with python-sc2 requirements, and successfully win most the combat scenarios. Meanwhile, the fine-tuned SMAC-R1 is able to infer correct scripts and successfully make decisions in the few-shot scenarios.
We validate the generated code within the StarCraft II client using 10 different random seeds. Both the times taken for coding are recorded as part of the results. For the sake of convenience, we use the open-source DeepSeek-V2.5 (deepSeek-Coder-v2.5-236B) as the base model for the planner, coder, and critic modules in API mode. However, these models can be replaced with other advanced alternatives, such as GPT-4.
@@ 4.1 Main Results
We present the results of win rates, coding rounds, coding rounds after SFT and DPO, and the final strategies planned by the LLM for each map in Table 1. As shown in the table, the LLM-generated code achieves near-perfect win rates on easier tasks and competitive results on more difficult tasks. Leveraging the LLM’s prior knowledge, the planner is able to select optimal strategies with relative ease, and thanks to the high coding capabilities of the DeepSeek model, most correct Python scripts are generated within 30 rounds of iteration.
@@
The planner LLM outlines the strategy framework by combining skills and defining the conditions under which they should be used. According to the results, "Focus Fire" emerges as one of the most common and effective skills across SMAC maps, as it quickly reduces the number of enemy units. For many of the easier tasks, once the planner identifies and selects the "Focus Fire" skill, the likelihood of generating a winning code script significantly increases. Conversely, on maps where focus fire is less effective, such as the corridor map, the LLM requires more rounds to produce alternative skills after receiving feedback from the rollout results.
@@
Figure 3: An example planner output that provides three potential skills: focus fire, hit and run, and Spread out and heal for the medivac unit. The planner also describes how to use each skill and when to use the skill.
Figure 4: An example coder output from the first round and after the critic promotion round given a main strategy skeleton. Both of the two example codes follow the strategy, but the latter code is more precise.
@@
Starting from a strategy skeleton, we present the coding result from the first round and the refined code following the promotion prompt from the critic module. Except for value adjustments and skill additions, the coder can enhance the implementation by accounting for more situational factors. As illustrated in Figure 4, in the kiting strategy, agents should retreat in the opposite direction to maintain a minimum distance from the closest enemy. However, on smaller maps, units may get stuck at the map’s edge, resulting in the retreating behavior being taken but failing. After several rounds of in-context learning improvements, the critic identifies that the planned strategy is sound, but the implementation lacks detail. Consequently, the coder begins dynamically calculating the retreat direction in the kiting method, improving the strategy’s overall effectiveness.
SMAC-R1 SFT and DPO
We fine-tune the Qwen2.5-7B-Base model by the scripts generated by the DeepSeek-coder-v2.5-236B model and their augmented scripts. We evaluate our model on the SMAC tasks and calculate the average coding times among 5 successful scripts which result in a higher than 90% winning rate. According to Table 1, the coding round with the critic module of DeepSeek is less than 30, but after the SFT and DPO rounds, the student model can generate correct codes in less than 5 rounds. Apart from that, we evaluate the student model on 10 newly designed but with similar features. According to Table 2, the student model can lock on the terrain advantage in the 3rp_vs_4zl map, the shield and healing management in 1mr1md_vs_1sc map, focus fire in 3mr_3mr and 5hd_vs_5hd scenarios and Split in 6mr_vs_1ac scenarios. This phenomenon also states the few-shot reasoning ability of the student LLM after the SFT and DPO.
Map Name WR % DeepSeek -236B Qwen2.5-7B (SFT+DPO) Planned Strategy
3m 100% 2 1.4 Focus Fire
8m 100% 5 1 Focus Fire
5m_vs_6m 0% - 1.4 Initial Forming, Focus Fire
8m_vs_9m 90% 22 2 Initial Forming, Focus Fire
10m_vs_11m 100% 19 1 Focus Fire, Grouping
25m 100% 32 1.2 Initial Forming, Focus Fire
27m_vs_30m 90% 35 1.2 Initial Forming, Focus Fire, Grouping
2s3z 100% 31 2 Focus Fire, Grouping
3s5z 100% 22 3 Focus Fire, Grouping
1c3s5z 100% 11 2.6 Engagement, Hit and Run, Flanking
3s5z_vs_3s6z 0% - - Focus Fire, Grouping, Kiting
2m_vs_1z 100% 4 1 Hit and Run
3s_vs_3z 100% 3 1 Focus Fire, Hit and Run
3s_vs_4z 100% 14 1.2 Focus Fire, Hit and Run
3s_vs_5z 100% 14 1.4 Focus Fire, Hit and Run
2c_vs_64zg 100% 18 1.2 Health and Shield Management, Terrain Advantage
corridor 100% 17 1 Health and Shield Management, Terrain Advantage
MMM 100% 15 3.6 Focus Fire, Hit and Run, Split out and Heal
MMM2 0% - - Focus Fire, Hit and Run, Split out and Heal
2s_vs_1sc 100% 23 1.2 Health and Shield Management, Terrain Advantage
bane_vs_bane 100% 14 2.4 Split
so_many_baneling 100% 11 1 Split
6h_vs_8z 0% - - Focus Fire, Grouping
Table 1: The winning rate over 10 times validation, the strategy planning rounds, the code generation rounds, and the planning results of each map in SMAC. We use the DeepSeek-Coder-V2.5 model with different parameters. The bugs are caused by the wrong API calls from the python-sc2 package in the generated codes. The results show the round number before the first successful generation for Deepseek and the average rounds of 5 successful generations for Qwen.
Agent Units Enemy Units Map Critical Strategy
1 Marauder and 1 Medivac 1 Spine Crawler Flat map Health Management, Heal
2 Void Ray 3 SporeCrawler Flat map Shield Management
3 Marauder 3 Marauder Flat map Focus Fire
3 Reaper 4 Zealot Cliff Terrain Advantage
5 Hydralisk 5 Hydralisk Flat map Initial Forming
6 Marine 1 Archon (Splash Damage) Flat map Split
6 Marine 1 SiegeTankSieged (Splash Damage) Flat map Split
3 Stalker 4 Zealot Cliff Kite, Path Finding
3 Hellion 24 Zergling Flat map Kite, Focus Fire
3 Reaper 24 Zergling Flat map Kite, Focus Fire
Table 2: New maps to test few-shot abilities of student LLM. Each map is similar to some original SMAC maps in the aspect of terrain features, shield features, or strategies.
SMAC-R1 GRPO optimization
Apart from applying iterative DPO to our SMAC-R1 model, we also integrate the GRPO algorithm using the reward calculation described earlier. The reward curves for different SMAC maps are presented in Figure 5 and Figure 6, while Table 3 provides the reward distribution statistics.
Figure 5: The learning curves among the 12/23 tasks in SMAC. The x-axis is the GRPO training steps and the y-axis is the average winning rate.(1/2)
Figure 6: The learning curves among the 11/23 tasks in SMAC. The x-axis is the GRPO training steps and the y-axis is the average winning rate.(2/2)
As shown in Figure 5 and Figure 6, the learning curves in most scenarios begin at a score of -1, indicating that the generated Python code is either missing from the response or incorrectly formatted. After limited training steps, the curves quickly approach -0.5, suggesting that SMAC-R1 can produce scripts, albeit with errors. Subsequently, the curves rapidly reach 0 and continue rising toward 1. It is worth noting that on the map 3s5z_vs_3s6z and the MMM2 which are super difficult in task configurations, our SMAC-R1 fails to generate correct scripts with the DPO process, yet generates scripts with certain winning rates after the GRPO processes. In the tasks such as 5m_vs_6m, it is hard for SMAC-R1 to explore an optimal strategy during the training process, so we believe that the model will generate an increasing number of correct samples with additional training steps.
Map Name Count Mean std Min Max Median
10m_vs_11m 3995 -0.045 0.483 -1.0 1.0 0.0
25m 3949 -0.028 0.489 -1.0 1.0 0.0
27m_vs_30m 3953 -0.193 0.367 -1.0 0.2 0.0
2c_vs_64zg 3968 -0.196 0.375 -1.0 0.5 0.0
2m_vs_1z 3946 -0.174 0.375 -1.0 1.0 0.0
2s3z 3994 -0.008 0.432 -1.0 1.0 0.0
2s_vs_1sc 3987 0.060 0.517 -1.0 1.0 0.0
3m 4000 -0.096 0.416 -1.0 1.0 0.0
3s5z 3984 0.470 0.681 -1.0 1.0 0.8
3s5z_vs_3s6z 3988 -0.144 0.340 -1.0 1.0 0.0
3s_vs_3z 3969 0.483 0.726 -1.0 1.0 1.0
3s_vs_4z 3968 0.349 0.702 -1.0 1.0 0.5
3s_vs_5z 3961 0.217 0.636 -1.0 1.0 0.1
5m_vs_6m 3980 -0.150 0.336 -1.0 0.0 0.0
6h_vs_8z 3969 -0.174 0.355 -1.0 0.0 0.0
8m 3966 0.406 0.740 -1.0 1.0 1.0
8m_vs_9m 4008 -0.135 0.393 -1.0 1.0 0.0
MMM 3953 0.203 0.732 -1.0 1.0 0.0
MMM2 3969 -0.243 0.368 -1.0 0.7 0.0
bane_vs_bane 3962 0.212 0.626 -1.0 1.0 0.4
corridor 4012 -0.131 0.391 -1.0 1.0 0.0
so_many_baneling 4019 -0.188 0.360 -1.0 0.3 0.0
Table 3: The number of testing instances and the GRPO training statistics, including the mean, std, min, max, and median value of each SMAC task.
5 Discussion
@@
Figure 8: An example indicates that the strategy learned from the 3s_vs_4z scenario is that one of the units is responsible for kiting, and another two units are responsible for attacking. Because the universal strategy is learned on the homogeneous scenario, the strategy is also suitable for 3, 4, and 5 enemy Zealots.
@@
Vanilla reinforcement learning models often suffer from low exploration efficiency, as they must train from scratch. Although some approaches in multi-agent reinforcement learning emphasize coordinated exploration techniques, these methods still explore at the action level, which limits their efficiency. In contrast, our SMAC-R1 framework leverages the prior knowledge of LLMs to explore at the skill level in the planner phase and at the decision-tree implementation level in the coder phase. This significantly enhances efficiency, allowing us to solve hard SMAC tasks in just tens of rounds of planning and code generation, whereas MARL algorithms typically require millions of timesteps over tens of thousands of episodes.
@@
Map descriptions provide crucial terrain advantages that enrich strategic planning. Some units, like the Colossus, can use the terrain to their advantage, such as climbing cliffs to kite enemies. Other units may use map edges to block choke points, reducing the risk of being surrounded by Zerglings. As illustrated in Figure 9, two Colossi can exploit the terrain by climbing up and down a cliff while most enemy Zerglings remain on the other side. Without this map information, strategies might unintentionally rely on map edge advantages. Another example is the corridor scenario, where units should hold choke points to prevent being surrounded by Zerglings. Without the choke point, the generated code is hard to win the corridor combat.
@@
Response Body Length
As shown in Figure 10, the average win rate exhibits a negative correlation with the length of the response body. The GRPO case study in the Appendix further supports this, indicating that longer preceding strategy analyses tend to produce more complex but less relevant implementations, ultimately leading to lower-quality code.
Figure 10: A comparison between the average winning rate and the length of response body during the GRPO training process.
Unlike DeepSeek-R1, where the response body progressively lengthens, SMAC-R1 generates increasingly shorter responses. This contrast highlights a fundamental difference between SMAC strategy reasoning and mathematical problem-solving. In math problem-solving, agents need to explore solutions through extensive reasoning. However, in SMAC decision-making, unnecessary strategies lead to wasted actions and a lower win rate. For instance, in the 3m scenario, agents must consistently focus fire on a single opponent until it is eliminated. If one agent retreats prematurely, the overall firepower decreases, increasing the likelihood of losing. As a result, the final strategy converges to the simplest yet most effective approach, avoiding unnecessary complexity.